2x-3y=6
4x+y=1
the question is solve the system of equations i believe this is where i can do elimination or substitution but im kind of confused
thank you in advance

Answer by scott8148(6628)

multiplying the 2nd equation by 3, and then adding it to the 1st equation ELIMINATES the y variable

(12x + 3y) + (2x - 3y) = 3 + 6 ___ 14x = 9 ___ x = 9/14


solving for y in the second equation, and then SUBSTITUTING the equivalent value of y into the first equation

y = 1 - 4x ___ 2x - 3(1 - 4x) = 6 ___ 2x - 3 + 12x = 6 ___ 14x = 9 ___ x = 9/14


both of these techniques finish by substituting the value of x into one of the equations to find the value of y